# UCM-Net: A Lightweight and Efficient Solution for Skin Lesion Segmentation using MLP and CNN

### v1

@@ 4.3 Ablation results
Figure 6: Stage Block Structures in ablation experiments.
@@
Furthermore, when incorporating the Group Loss into the UCM-Net architecture, denoted as "UCM-Net + Group Loss", the model’s performance continued to excel. This enhancement resulted in a higher mIoU of 80.63% and a mDice of 87.64%, demonstrating the effectiveness of the proposed Group Loss in further improving segmentation accuracy. Figures 7-8 show that "UCM-Net + Group Loss" always presents the high scores of mIoU and mDice with the training epoch increase.
The above findings in the ablation experiments underscore the significance of architectural innovations such as the UCMNet block in achieving superior semantic segmentation performance, even with fewer parameters and computational complexity than the U-Net baseline.
@@ 5 Conclusion
This paper introduces UCM-Net, a novel, lightweight, and highly efficient solution. UCM-Net combines MLP and CNN, providing robust feature learning capabilities while maintaining a minimal parameter count and reduced computational demand. We applied this innovative approach to the challenging task of skin lesion segmentation, conducting comprehensive experiments with a range of evaluation metrics to showcase its effectiveness and efficiency. The results of our extensive experiments unequivocally demonstrate UCM-Net’s superior performance compared to the state-of-the-art EGE-UNet. Remarkably, UCM-Net is the first model with fewer than 50KB parameters and consuming less than 0.05 GLOPs for skin lesion segmentation. Looking forward to future research endeavors, we aim to expand the application of UCM-Net to other critical medical image tasks, advance the field, and explore how this efficient architecture can contribute to a broader spectrum of healthcare applications, potentially revolutionizing how we utilize deep learning for medical image analysis.

### v4

@@ 4.3 Ablation results
Furthermore, when incorporating the proposed Group Loss into the UCM-Net architecture, denoted as "UCM-Net + Group Loss(ours)", the model’s performance continued to excel. This enhancement resulted in a higher mIoU of 81.00% and a mDice of 87.98%, demonstrating the effectiveness of the proposed Group Loss in further improving segmentation accuracy. Figures 7-8 show that "UCM-Net + Group Loss(ours)" always presents the high scores of mIoU and mDice with the training epoch increase.
The above findings in the ablation experiments underscore the significance of architectural innovations such as the UCMNet block in achieving superior semantic segmentation performance, even with fewer parameters and computational complexity than the U-Net baseline. We note the batch size will affect the prediction performance. So, we tested our model with different batch sizes and reported it in Table 6 in the appendix.
@@ 5 Conclusion
This paper introduces UCM-Net, a novel, lightweight, and highly efficient solution. UCM-Net combines MLP and CNN, providing robust feature learning capabilities while maintaining a minimal parameter count and reduced computational demand. We applied this innovative approach to the challenging task of skin lesion segmentation, conducting comprehensive experiments with a range of evaluation metrics to showcase its effectiveness and efficiency. The results of our extensive experiments unequivocally demonstrate UCM-Net’s superior performance compared to the state-of-the-art EGE-UNet. Remarkably, UCM-Net is the first model with fewer than 50KB parameters and consuming less than 0.05 GLOPs for skin lesion segmentation. Looking forward to future research endeavors, we aim to expand the application of UCM-Net to other critical medical image tasks, advancing the field and exploring how this efficient architecture can contribute to a broader spectrum of healthcare applications. This potential revolution in utilizing deep learning for medical image analysis opens up numerous possibilities for enhancing patient care and diagnostic accuracy. However, it is essential to acknowledge that compared to models leveraging pre-trained components, UCM-Net currently demonstrates lower accuracy performance. Recognizing this limitation, our future efforts will focus on optimizing UCM-Net to improve its accuracy. Such advancements are crucial for ensuring that the model maintains its efficiency and competes favorably in performance with existing state-of-the-art solutions. By addressing these challenges, we aim to advance the field further and expand the impact of deep learning in healthcare applications, making significant contributions to medical imaging and beyond.